OpenStudy (anonymous):

Rationalize denominator: 5/8+sqrt (x)

OpenStudy (anonymous):

(5+8sqrt(x))/8

OpenStudy (anonymous):

not it...

OpenStudy (anonymous):

\[\frac{5}{8+\sqrt{x}}=\frac{5(8-\sqrt{x})}{64-x}\]

OpenStudy (anonymous):

sqrt(5)/sqrt(8) = sqrt(40)/8 = 2*sqrt(10)/8 = sqrt(10)/4